Newton's law of cooling states that the temperature of an object changes at a rate proportional to the difference between its temperature and that of its surroundings. Suppose that the temperature of a cup of coffee obeys Newton's law of cooling. If the coffee has a temperature of 200 degrees Fahrenheit when freshly poured, and 3 minutes later has cooled to 180 degrees in a room at 76 degrees, determine when the coffee reaches a temperature of 150 degrees.
The coffee will reach a temperature of 150 degrees in ______ minutes.
Answer #1

Please see the attached images..

ЯН) Мол TIMELIT 2. 12 According to Newtons how of cooling: d 1 - K (T-To) - 0 s temp of Coffee Is surrounding temp. T(0) = 2104 - ek 124 en 10 w = -3x = -1 en Holly – 0.0586 Tlt) = 124e -0.05864 + 76 = 1240-0.0586t. 150 = 12re + 76 -0.05867 – 150-76